OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

Since parallel means the same slope, use y = 3x + b y = 3x + b Plug in the coordinate (3, -2) and solve for b. -2 = 3(3) + b

OpenStudy (anonymous):

uhm, is it y = -3x - 12 ?

OpenStudy (anonymous):

Um no. -2 = 9 + b Canb you finish solving for b?